OpenStudy (anonymous):

Soybean meal is 12% protein ; cornmeal is 6%. How many pounds of each should be mixed together in order to get 240-Ib mixture that is 9% protein ? How many pounds of the corn meal should be in the mixture ? _____ pounds How many pounds of the soybean meal should be in the mixture ______ pounds

OpenStudy (saifoo.khan):

cornmeal: 120lb soyabean: 120lb

OpenStudy (anonymous):

saifoo you sure

OpenStudy (saifoo.khan):

Yes, 101%. ^_^

OpenStudy (saifoo.khan):

Let me show you.

OpenStudy (saifoo.khan):

Since 9% is halfway between 12% and 6%, it would need 120 lb of each However, here is the algebra procedure: Let x = amt of soybean meal (12%) It says the resulting amt = 240 lb, therefore: (240-x) = amt of cornmeal (6%) write the decimal equiv of the per cent protein equation: 0.12x + 0.06(240-x) = 0.09(240) 0.12x + 14.4 - 0.06x = 21.6 0.12x - 0.06x = 21.6 - 14.4 0.06x = 7.2 x = 7.2/0.06 x = 120.